In 2018, Lyra Co. acquired land containing ore for a total cost of ₱60,000,000 immovable tangible equipment costs totaled ₱5,000,000. Geologist estimate that the total units estimated to be extracted is ₱2,000,000. It is estimated that 200,000 units will be extracted each year during the useful life of the wasting assets. The immovable equipment has a useful life of 15 years.

Actual units extracted during the years 2018 to 2020 totaled 620,000 units. There were no units extracted during 2021 because of employee strike.

In 2022, the company resumed operations and there were 150,000 units extracted.

Questions:

Based on the above data, answer the following:

1. How much is the 2021 depreciation of the immovable equipment?

2. How much is the 2022 depreciation of the immovable equipment?



1
Expert's answer
2021-06-16T13:32:30-0400

Solution:

1.). The 2021 depreciation of the immovable equipment:

Calculation of Depreciation using Units of Production method:

Annual depreciation ="Depreciable\; value\times (\frac{Actual\; units \; produced \; during\; the \; year}{Estimated\; total \; production\; over\; the\; assets\; useful\; life})"


Depreciable value = Original value – Salvage value

Original value = 5,000,000

Salvage value = 0

Depreciable value = 5,000,000 – 0 = 5,000,000

Actual units produced during 2021 = 0

Estimated production over the asset’s useful life = 200,000 "\times" 15 = 3,000,000


Annual depreciation = "5,000,000\times (\frac{0}{3,000,000} ) = 0"

The 2021 depreciation of the immovable equipment = ₱0

 

2.). The 2022 depreciation of the immovable equipment:

Annual depreciation = Depreciable value x (Actual units produced during the year/Estimated total production over the asset useful life)

Depreciable value = Original value – Salvage value

Original value = 5,000,000

Salvage value = 0

Depreciable value = 5,000,000 – 0 = 5,000,000

Actual units produced during 2021 = 150,000

Estimated production over the asset’s useful life = 200,000 "\times"15 = 3,000,000


Annual depreciation = "5,000,000\times (\frac{150,000}{3,000,000} ) = 250,000"


The 2021 depreciation of the immovable equipment = ₱250,000